# Quillen’s conjecture and unitary groups

### v3

@@ 1. Introduction
In [5, Corollary 2], K.S. Brown proved the following result: If $\mathcal{S}_{p}(G)$ is the poset of the non-trivial $p$-subgroups of a finite group $G$ ordered by inclusion, and $|\mathcal{S}_{p}(G)|$ denotes its topological realization, then $\chi(|\mathcal{S}_{p}(G)|)\equiv 1$ modulo the highest power of $p$ dividing the order of $G$, where $\chi$ stands for Euler characteristic. Thus, here, the poset $\mathcal{S}_{p}(G)$ relates a topological invariant to an algebraic invariant and, since then, $p$-subgroups posets have linked algebraic topology and finite group theory from other perspectives, including homotopy equivalences and Tits buildings [19], group cohomology and geometries [22], homology decompositions [9, 4], and, more recently, fusion systems [1].
@@
The greatest advances towards proving the conjecture in full generality are [3] by Aschbacher and Smith and [18] by Piterman and Smith. These works are based on the analysis of the components of a hypothetical minimal counterexample to the conjecture, and they employ the CFSG. Very roughly, the method in [3] states that, under certain group-theoretic reductions on a group $G$ and for $p>5$, the following dichotomy should hold: for any component $L$ of $G$, either every $p$-extension of $L$ satisfies the Quillen dimension property for $p$, or $L$ possesses a so-called Robinson subgroup (see [3, p.491] and [20]).
In order to explain the notions employed in the previous paragraph, let $\operatorname{m}_{p}(G)$ denote the $p$-rank of the finite group $G$, and recall that this number is the largest integer $m$ such that $G$ has an elementary abelian $p$-subgroup of order $p^{m}$. Then we say that $G$ has the Quillen-dimension property for $p$, written as $\mathcal{Q}\mathcal{D}_{p}$, if $|\mathcal{A}_{p}(G)|$ has non-zero reduced rational homology in its maximal dimension, i.e., in $\operatorname{m}_{p}(G)-1$, see Definition 3.10. This property was implicitly employed by Quillen to establish his conjecture in the aforementioned cases, and plays a central role in the works of Aschbacher and Smith and of Piterman and Smith. On the other hand, the $p$-extensions of a group $L$ are the split extensions $LB$ of $L$ by an elementary abelian $p$-group $B$ of outer automorphisms of $L$. The case $B=1$ is allowed, so that $L$ is a $p$-extension of itself. The components of a group are defined in Section 2.
@@
Theorem ([3, Main Theorem]).
@@
In [6] and [7], a new geometric method is introduced and employed to prove the Quillen dimension property in known cases, namely, for $p$-solvable groups and for groups of type alternating, symmetric, linear, unitary, symplectic, and orthogonal. This approach works but for a few exceptions, including the case $\operatorname{PSU}\nolimits_{n}(q)$ when $p$ is odd and $p\mid q+1$. In the current work, we elaborate on this geometric method and show that unitary groups and their $p$-extensions do satisfy the Quillen dimension property, as conjectured by Aschbacher and Smith in [3, Conjecture 4.1], see Theorem B below. We proceed via a seemingly indirect path, as we deal first with general unitary groups.
@@
This theorem shows that [3, Conjecture 4.1] is valid. Moreover, by the result [3, Main Theorem] by Aschbacher and Smith and its extension to all odd primes in [18, Theorem 1.1] by Piterman and Smith, we conclude the following result.
@@
There are several reasons that explain why these techniques do not cover yet the case $p=2$ . First, Theorems A and B and most of the aforementioned works that establish the Quillen dimension property for odd $p$ in various cases do not immediately extend to $p=2$. Some obstructions to handle are that the description of $p$-ranks, maximal elementary abelian $p$-subgroups, and outer automorphism groups, are more involved for $p=2$ than for odd $p$, see [3, Section 2]. Secondly, a dichotomy similar to the one introduced above for odd $p$ is not yet completed for $p=2$, in particular, a so-called $\mathcal{Q}\mathcal{D}$-list of simple groups possibly failing the Quillen dimension property, cf. [3, Theorem 3.1], is still lacking. As mentioned above, relevant progress in this direction has been made by Piterman and Smith in [18], and other elimination results, besides the Quillen dimension property, are needed to prepare this missing list.
@@ 2. Group theoretical preliminaries
    a cyclic group prime order,
@@
    one of twenty-six sporadic groups.
Only the groups in (i) are commutative, so that the groups in (ii,iii,iv) are the non-abelian simple groups. The projective special unitary group $\operatorname{PSU}\nolimits_{n}(q)$, which is defined below and is the theme of this paper, is a finite simple group of Lie type.
The simple groups are the building bricks of any finite group and, with the aim of finding out the simple groups involved in $G$, we say that $H\leq G$ is subnormal if there is a subnormal series starting in $H$ and ending in $G$, i.e., a chain of subgroups such that every subgroup of the chain is normal in the next one,
$H=H_{1}\lhd H_{2}\lhd\ldots\lhd H_{n-1}\lhd H_{n}=G.$
If the subnormal series cannot be refined further, i.e., if $H=1$ and every subgroup is maximal in the next one, the chain is called a composition series. Then the quotients groups $H_{i}/H_{i-1}$ are simple, are called the composition factors of $G$, and constitute the simple groups involved in $G$. For instance, if all these quotients are abelian, then the group $G$ is termed solvable. In the opposite direction, a subgroup $H\leq G$ is a component of $G$ if it is subnormal and, in addition, it is quasi-simple, i.e., it satisfies that $H=[H,H]$ and that $H/Z(H)$ is simple. The special unitary group $\operatorname{SU}\nolimits_{n}(q)$ is, in general, quasi-simple, see below. If $G$ has a non-trivial component, then its simple quotient is a non-abelian composition factor of $G$ and, in particular, $G$ is not solvable. As discussed in the introduction, the analysis of the components of given finite group is a cornerstone for the results of Aschbacher and Smith [3] and of Piterman and Smith [18].
@@
Throughout the paper we assume $n\geq 2$, as for $n=1$ we would get abelian groups in the forthcoming constructions. Let $\operatorname{GL}\nolimits_{n}(q)$ denote the general linear group over the field $\mathbb{F}_{q}$ of $q=s^{l}$ elements, $s$ a prime, and let $\operatorname{SL}\nolimits_{n}(q)$ denote the special linear group, which is the normal subgroup of $\operatorname{GL}\nolimits_{n}(q)$ consisting of the matrices of determinant $1$. In order to easily describe the matrices of $\operatorname{GL}\nolimits_{n}(q)$ appearing in the rest of this work, see Subsection 2.3 for instance, we introduce the notation
(1) $\Psi_{l_{1},\ldots,l_{k}}(x,y),$
where $x=(x_{ij})\in\operatorname{GL}\nolimits_{k}(q)$ and $y=(y_{ij})\in\operatorname{GL}\nolimits_{n}(q)$, to denote the matrix of $\operatorname{GL}\nolimits_{n}(q)$ equal to $y$ but for substituting $y_{l_{i}l_{j}}$ by $x_{ij}$. We also denote by $I_{n}$ the $n\times n$ identity matrix.
Recall now that the field $\mathbb{F}_{q^{2}}$ of $q^{2}=s^{2l}$ elements has the involution
@@
and that $\mathbb{F}_{q}=\{\alpha\in\mathbb{F}_{q^{2}}|\alpha=\overline{\alpha}\}$. Over an $n$-dimensional vector space over $\mathbb{F}_{q^{2}}$, we may define the following sesquilinear form employing the involution above,
@@
We define the general unitary group $\operatorname{GU}\nolimits_{n}(q)$ as the subgroup of orthonormal matrices of $\operatorname{GL}\nolimits_{n}(q^{2})$, i.e.,
@@
where $\overline{x}$ is obtained from $x$ by applying $\overline{\cdot}$ to each entry and $^{\operatorname{T}\nolimits}$ denotes transpose. Analogously, we define the special unitary group $\operatorname{SU}\nolimits_{n}(q)$ as the subgroup of orthonormal matrices of $\operatorname{SL}\nolimits_{n}(q^{2})$.
@@
have $O_{p}(G)\neq 1$ for $p=2,3$ the only prime dividing $q+1$, and the same happens to the groups $\operatorname{PGU}\nolimits_{2}(2)$, $\operatorname{PGU}\nolimits_{2}(3)$, and $\operatorname{PGU}\nolimits_{3}(2)$, and to the extensions of $\operatorname{PSU}\nolimits_{3}(8)$ and of $\operatorname{PGU}\nolimits_{3}(8)$ by field automorphisms of order $3$. Thus, by “conical contractibility” [19, Proposition 2.4], $\mathcal{Q}\mathcal{D}_{p}$ cannot hold for these groups, explaining the cases excluded from Theorems A, B, and C. To finish this subsection, note that, $\operatorname{PGU}\nolimits_{n}(q)$ is not simple as it contains $\operatorname{PSU}\nolimits_{n}(q)$ as a normal subgroup,
@@
see [12, §2.7] and [10, I.7] for more details. All automorphisms of $\operatorname{PSU}\nolimits_{n}(q)$ are induced by automorphisms of $\overline{G}$ that commute with $\sigma$, see [12, Theorem 2.5.4] for a precise statement, and those automorphisms of $\overline{G}$ induce also automorphisms of $\operatorname{PGU}\nolimits_{n}(q)$. Below, we define, following [12, Definition 2.5.13], the inner-diagonal and field automorphisms of $\operatorname{PSU}\nolimits_{n}(q)$, and we will employ the same names to denote the corresponding automorphisms induced in $\operatorname{PGU}\nolimits_{n}(q)$.
Thus, we declare the inner-diagonal automorphisms of $\operatorname{PSU}\nolimits_{n}(q)$ (and of $\operatorname{PGU}\nolimits_{n}(q)$) as those induced by conjugation by elements from $\operatorname{PGU}\nolimits_{n}(q)$. In addition, the Frobenius map induces the automorphism of $\operatorname{SU}\nolimits_{n}(q)$ (and of $\operatorname{GU}\nolimits_{n}(q)$) that acts as $\lambda\mapsto\lambda^{s}$ on the matrix entries, and descends to an automorphism $\phi$ of $\operatorname{PSU}\nolimits_{n}(q)$ (and of $\operatorname{PGU}\nolimits_{n}(q)$). We define, see [12, Definition 2.5.13(c1)], the field automorphisms of $\operatorname{PSU}\nolimits_{n}(q)$ (and of $\operatorname{PGU}\nolimits_{n}(q)$) as the $\operatorname{PGU}\nolimits_{n}(q)$-conjugates of a power of $\phi$ of odd order. In particular, with these definitions, all cyclic subgroups of field automorphisms of $\operatorname{PGU}\nolimits_{n}(q)$ of the same odd order are $\operatorname{PGU}\nolimits_{n}(q)$-conjugated. We are interested in the case that there exists a field automorphism of odd order $p$.
@@
The values for the $p$-ranks in the display of the statement are obtained from [10, I.10-6(1)] and [10, I.10-2(2)]. For the $p$-extensions, recall that, for $n\geq 3$, we have, by [8, Chapitre IV, §7], that
@@
but for $\operatorname{PSU}\nolimits_{3}(2)$, which is included in the non-simple exceptions of Equation (2). Quotening by the inner automorphisms, we get the outer automorphisms,
@@
where $C_{(n,q+1)}$ is isomorphic the quotient group in Equation (3), $C_{2l}$ is generated by $\phi$, and the action for this semi-direct product is described in [12, Theorem 2.5.12(g)]. This gives rise to the types of $p$-extensions in points (i-iv) of the statement. (We note that the full automorphism group of $\operatorname{PGU}\nolimits_{n}(q)$ is only fully understood for $n=5$ or $n\geq 7$, see more details in [8, Chapitre IV, §7].) Finally, for the $p$-ranks of the $p$-extensions, we use that $\operatorname{PSU}\nolimits_{n}(q)\subseteq\operatorname{PGU}\nolimits_{n}(q)$. ∎
@@
For $f$ the sesquilinear form introduced in Subsection 2.1, we say that a vector $v$ is isotropic if $f(v,v)=0$ and that it is anisotropic otherwise. Then, an anisotropic vector $v$ and an element $\alpha\in\mathbb{F}^{*}_{q^{2}}$ with $\alpha\overline{\alpha}=1$, determine an element of $\operatorname{GU}\nolimits_{n}(q)$, called quasi-reflection, as follows,
@@
A simple computation shows the following properties, all of which are well-known,
(4a) $X_{\gamma v,\alpha}$ $=X_{v,\alpha}\text{ for all $\gamma\in\mathbb{F}^{*}_{q^{2}}$,}$
(4b) $X_{v,\alpha}X_{v,\alpha^{\prime}}$ $=X_{v,\alpha\alpha^{\prime}},$
(4c) $X_{v,\alpha}X_{v^{\prime},\alpha^{\prime}}$ $=X_{v^{\prime},\alpha^{\prime}}X_{v,\alpha}\Leftrightarrow f(v,v^{\prime})=0\text{ or $v^{\prime}=\gamma v$ with $\gamma\in\mathbb{F}_{q^{2}}$,}$
(4d) ${}^{{x}\!}{X}_{v,\alpha}$ $=X_{x(v),\alpha},\text{ for any element $x\in\operatorname{GU}\nolimits_{n}(q)$,}$
(4e) $X_{v,\alpha}(v)$ $=\alpha v\text{, }X_{v,\alpha}(u)=u\text{ if $f(u,v)=0$.}$
@@
If $n\geq 2$ and $q\geq 3$, then there exists a non-diagonal quasi-reflection $x\in\operatorname{GU}\nolimits_{n}(q)$, such that, for any diagonal element $t=\operatorname{diag}(\alpha_{1},\ldots,\alpha_{n})\in T$, we have that $^{{x}\!}{t}$ is non-diagonal if $\alpha_{1}\neq\alpha_{2}$ and that $^{{x}\!}{t}=t$ if $\alpha_{1}=\alpha_{2}$.
@@
For arbitrary values $\beta,\gamma\in\mathbb{F}^{*}_{q^{2}}$ satisfying Equation (5), a straightforward computation shows the following, where we use the notation of Equation (1),
(6) $X_{v,\gamma}=\Psi_{1,2}(X_{(1,\beta)^{\operatorname{T}\nolimits},\gamma},I_{n})$
with
$X_{(1,\beta)^{\operatorname{T}\nolimits},\gamma}=\frac{1}{\zeta}\begin{pmatrix}\gamma+\beta\overline{\beta}&(\gamma-1)\overline{\beta}\\ (\gamma-1)\beta&1+\gamma\beta\overline{\beta}\end{pmatrix}$
and $\zeta=1+\beta\overline{\beta}$, as well as that
$^{{X_{v,\gamma}}\!}{t}=\Psi_{1,2}({}^{{\left(X_{(1,\beta)^{\operatorname{T}\nolimits},\gamma}\right)}\!}{\operatorname{diag}}(\theta_{1},\theta_{2}),t),$
[…]
$\frac{1}{\zeta^{2}}\begin{pmatrix}\alpha_{1}\zeta^{2}+(\alpha_{2}-\alpha_{1})(\gamma-1)(\overline{\gamma}-1)\beta\overline{\beta}&(\alpha_{1}-\alpha_{2})(\overline{\gamma}-1)\overline{\beta}(\gamma+\beta\overline{\beta})\\ (\alpha_{1}-\alpha_{2})(\gamma-1)\beta(\overline{\gamma}+\beta\overline{\beta})&\alpha_{2}\zeta^{2}+(\alpha_{1}-\alpha_{2})(\gamma-1)(\overline{\gamma}-1)\beta\overline{\beta}\end{pmatrix}$,
Thus, for the conclusion of the proposition to hold, we need to choose $\beta,\gamma\in\mathbb{F}_{q^{2}}$ satisfying Equation (5) together with

### v5

@@ 1. Introduction
In [6, Corollary 2], K.S. Brown proved the following result: If $\mathcal{S}_{p}(G)$ is the poset of the non-trivial $p$-subgroups of a finite group $G$ ordered by inclusion, and $|\mathcal{S}_{p}(G)|$ denotes its topological realization, then $\chi(|\mathcal{S}_{p}(G)|)\equiv 1$ modulo the highest power of $p$ dividing the order of $G$, where $\chi$ stands for Euler characteristic. Thus, here, the poset $\mathcal{S}_{p}(G)$ relates a topological invariant to an algebraic invariant and, since then, $p$-subgroups posets have linked algebraic topology and finite group theory from other perspectives, including homotopy equivalences and Tits buildings [19], group cohomology and geometries [22], homology decompositions [5, 9], and, more recently, fusion systems [2].
@@
The greatest advances towards proving the conjecture in full generality are [4] by Aschbacher and Smith and [18] by Piterman and Smith. These works are based on the analysis of the components of a hypothetical minimal counterexample to the conjecture, and they employ the CFSG. Very roughly, the method in [4] states that, under certain group-theoretic reductions on a group $G$ and for $p>5$, the following dichotomy should hold: for any component $L$ of $G$, either every $p$-extension of $L$ satisfies the Quillen dimension property for $p$, or $L$ possesses a so-called Robinson subgroup (see [4, p.491] and [20]).
In order to explain the notions employed in the previous paragraph, let $\operatorname{m}_{p}(G)$ denote the $p$-rank of the finite group $G$, and recall that this number is the largest integer $m$ such that $G$ has an elementary abelian $p$-subgroup of order $p^{m}$. Then we say that $G$ has the Quillen-dimension property for $p$, written as $\mathcal{Q}\mathcal{D}_{p}$, if $|\mathcal{A}_{p}(G)|$ has non-zero reduced rational homology in its maximal dimension, i.e., in degree $\operatorname{m}_{p}(G)-1$, see Definition 3.10. This property was implicitly employed by Quillen to establish his conjecture in the aforementioned cases, and plays a central role in the works of Aschbacher and Smith and of Piterman and Smith. On the other hand, the $p$-extensions of a group $L$ are the split extensions $LB$ of $L$ by an elementary abelian $p$-group $B$ of outer automorphisms of $L$. The case $B=1$ is allowed, so that $L$ is a $p$-extension of itself. The components of a group are defined in Section 2.
@@
Theorem ([4, Main Theorem]).
@@
In [7] and [8], a new geometric method is introduced and employed to prove the Quillen dimension property for $p$ in known cases, namely, for $p$-solvable groups if $p$ is any prime, and for groups of type alternating, symmetric, linear, unitary, symplectic, and orthogonal for some odd primes $p$. This approach works but for a few exceptions, including the case $\operatorname{PSU}\nolimits_{n}(q)$ when $p$ is odd and $p\mid q+1$. In the current work, we elaborate on this geometric method and show that unitary groups and their $p$-extensions do satisfy the Quillen dimension property, as conjectured by Aschbacher and Smith in [4, Conjecture 4.1], see Theorem B below. We proceed via a seemingly indirect path, as we deal first with general unitary groups.
@@
This theorem shows that [4, Conjecture 4.1] is valid. Moreover, by the result [4, Main Theorem] by Aschbacher and Smith and its extension to all odd primes in [18, Theorem 1.1] by Piterman and Smith, we conclude the following result.
@@
There are several reasons that explain why these techniques do not cover yet the case $p=2$ . First, Theorems A and B and most of the aforementioned works that establish the Quillen dimension property for odd $p$ in various cases do not immediately extend to $p=2$. Some obstructions to handle are that the description of $p$-ranks, maximal elementary abelian $p$-subgroups, and outer automorphism groups, are more involved for $p=2$ than for odd $p$, see [4, Section 2]. Secondly, a dichotomy similar to the one introduced above for odd $p$ is not yet completed for $p=2$, in particular, a so-called $\mathcal{Q}\mathcal{D}$-list of simple groups possibly failing the Quillen dimension property, cf. [4, Theorem 3.1], is still lacking. As mentioned above, relevant progress in this direction has been made by Piterman and Smith in [18], and other elimination results, besides the Quillen dimension property, are needed to prepare this missing list.
@@ 2. Group theoretical preliminaries
    a cyclic group of prime order,
@@
    one of the twenty-six sporadic groups.
Only the groups in (i) are abelian, so that the groups in (ii,iii,iv) are the non-abelian simple groups. The projective special unitary group $\operatorname{PSU}\nolimits_{n}(q)$, which is defined below and is the theme of this paper, is a finite group of Lie type, and it is simple for $n\geq 2$ except for a few exceptions for $n=2$, see below.
The components of a finite group are defined as its quasi-simple subnormal subgroups, and we recall that a finite group $H$ is quasi-simple if it satisfies that $H=[H,H]$ and that $H/Z(H)$ is simple. The special unitary group $\operatorname{SU}\nolimits_{n}(q)$ is, in general, quasi-simple, see below. As discussed in the introduction, the analysis of the components of given finite group is a cornerstone for the results of Aschbacher and Smith [4] and of Piterman and Smith [18].
@@
Throughout the paper we assume $n\geq 2$, as for $n=1$ we would get abelian groups in the forthcoming constructions. Let $\operatorname{GL}\nolimits_{n}(q)$ denote the general linear group over the field $\mathbb{F}_{q}$ of $q=s^{l}$ elements, $s$ a prime, and let $\operatorname{SL}\nolimits_{n}(q)$ denote the special linear group, which is the normal subgroup of $\operatorname{GL}\nolimits_{n}(q)$ consisting of the matrices of determinant $1$.
We denote by $I_{n}$ the $n\times n$ identity matrix, and recall that the field $\mathbb{F}_{q^{2}}$ of $q^{2}=s^{2l}$ elements has the involution automorphism
@@
and that $\mathbb{F}_{q}=\{\alpha\in\mathbb{F}_{q^{2}}|\alpha=\overline{\alpha}\}$. Over an $n$-dimensional vector space over $\mathbb{F}_{q^{2}}$, we may define the following Hermitian form employing the involution above,
@@
We define the general unitary group $\operatorname{GU}\nolimits_{n}(q)$ as the subgroup of unitary matrices of $\operatorname{GL}\nolimits_{n}(q^{2})$ with respect to the form $f$, i.e.,
@@
where $\overline{x}$ is obtained from $x$ by applying $\overline{\cdot}$ to each entry and $^{\operatorname{T}\nolimits}$ denotes transpose. Analogously, we define the special unitary group $\operatorname{SU}\nolimits_{n}(q)$ as the subgroup of unitary matrices of $\operatorname{SL}\nolimits_{n}(q^{2})$.
@@
have $O_{p}(G)\neq 1$ for $p=3,2,3$ respectively (note this is the only prime dividing $q+1$), and the same happens to the groups $\operatorname{PGU}\nolimits_{2}(2)$, $\operatorname{PGU}\nolimits_{2}(3)$, and $\operatorname{PGU}\nolimits_{3}(2)$, and to the extensions of $\operatorname{PSU}\nolimits_{3}(8)$ and of $\operatorname{PGU}\nolimits_{3}(8)$ by field automorphisms of order $3$ with $p=3$. Thus, by “conical contractibility” [19, Proposition 2.4], $\mathcal{Q}\mathcal{D}_{p}$ cannot hold for these groups, explaining the cases excluded from Theorems A and B. To finish this subsection, we recall the description of the quotient of $\operatorname{PGU}\nolimits_{n}(q)$ by $\operatorname{PSU}\nolimits_{n}(q)$,
@@
see [12, §2.7] and [10, I.7] for more details. All automorphisms of $\operatorname{PSU}\nolimits_{n}(q)$ are induced by automorphisms of $\overline{G}$ that commute with $\sigma$, see [12, Theorem 2.5.4] for a precise statement, and such automorphisms also induce automorphisms of $\operatorname{PGU}\nolimits_{n}(q)$. Below, we define, following [12, Definition 2.5.13], the diagonal, inner-diagonal, and field automorphisms of $\operatorname{PSU}\nolimits_{n}(q)$, and we will employ the same names to denote the corresponding automorphisms induced in $\operatorname{PGU}\nolimits_{n}(q)$.
Thus, we declare the non-trivial diagonal automorphisms of $\operatorname{PSU}\nolimits_{n}(q)$ (and of $\operatorname{PGU}\nolimits_{n}(q)$) as those induced by conjugation by elements from $\operatorname{PGU}\nolimits_{n}(q)\setminus\operatorname{PSU}\nolimits_{n}(q)$, and the inner-diagonal automorphisms of $\operatorname{PSU}\nolimits_{n}(q)$ (and of $\operatorname{PGU}\nolimits_{n}(q)$) as those induced by conjugation by elements from $\operatorname{PGU}\nolimits_{n}(q)$. In addition, the Frobenius map induces the automorphism of $\operatorname{SU}\nolimits_{n}(q)$ (and of $\operatorname{GU}\nolimits_{n}(q)$) that acts as $\lambda\mapsto\lambda^{s}$ on the matrix entries, and descends to an automorphism $\phi$ of $\operatorname{PSU}\nolimits_{n}(q)$ (and of $\operatorname{PGU}\nolimits_{n}(q)$). We define, see [12, Definition 2.5.13(c1)], the field automorphisms of $\operatorname{PSU}\nolimits_{n}(q)$ (and of $\operatorname{PGU}\nolimits_{n}(q)$) as the $\operatorname{PGU}\nolimits_{n}(q)$-conjugates of a power of $\phi$ of odd order. In particular, with these definitions, all cyclic subgroups of field automorphisms of $\operatorname{PGU}\nolimits_{n}(q)$ of the same odd order are $\operatorname{PGU}\nolimits_{n}(q)$-conjugated. We are interested in the case that there exists a field automorphism of odd order $p$.
@@
The values for the $p$-ranks in the display of the statement are obtained from [10, I.10-6(1)] and [10, I.10-2(2)]. For the $p$-extensions, recall that, for $n\geq 3$, we have, by [12, Theorem 2.5.12], that
@@
but for $\operatorname{PSU}\nolimits_{3}(2)$, which is included in the non-simple exceptions of Equation (1). Quotienting by the inner automorphism group, we get the outer automorphism group,
@@
where $C_{(n,q+1)}$ is isomorphic the quotient group in Equation (2), $C_{2l}$ is generated by $\phi$, and the action for this semi-direct product is described in [12, Theorem 2.5.12(g)]. This gives rise to the types of $p$-extensions in points (i-iv) of the statement by [12, Proposition 4.9.1(d)]. Finally, for the $p$-ranks of the $p$-extensions, we use that $\operatorname{PSU}\nolimits_{n}(q)\subseteq\operatorname{PGU}\nolimits_{n}(q)$. ∎
@@
For $f$ the Hermitian form introduced in Subsection 2.1, we say that a vector $v$ is isotropic if $f(v,v)=0$ and that it is anisotropic otherwise. Then, an anisotropic vector $v$ and an element $\alpha\in\mathbb{F}^{*}_{q^{2}}$ with $\alpha\overline{\alpha}=1$, determine an element of $\operatorname{GU}\nolimits_{n}(q)$, called quasi-reflection, as follows, see [13, p.94],
@@
If $n\geq 2$ and $q\geq 3$, then there exists a quasi-reflection $x\in\operatorname{GU}\nolimits_{n}(q)$, such that $x$ is not a monomial matrix and that, for any diagonal element $t=\operatorname{diag}(\alpha_{1},\ldots,\alpha_{n})\in T$, $^{{x}\!}{t}$ is non-diagonal if $\alpha_{1}\neq\alpha_{2}$ and $^{{x}\!}{t}=t$ if $\alpha_{1}=\alpha_{2}$.
@@
It is straightforward that $x$ defined in this way is the block diagonal matrix
(4) $x=\operatorname{diag}(X_{(1,\beta)^{\operatorname{T}\nolimits},\gamma},I_{n-2}).$
In particular, it is enough to solve the case $n=2$, and we do so below.
For arbitrary values $\beta,\gamma\in\mathbb{F}^{*}_{q^{2}}$ satisfying Equation (3), a straightforward computation shows that
$X_{(1,\beta)^{\operatorname{T}\nolimits},\gamma}=\frac{1}{\zeta}\begin{pmatrix}\gamma+\beta\overline{\beta}&(\gamma-1)\overline{\beta}\\ (\gamma-1)\beta&1+\gamma\beta\overline{\beta}\end{pmatrix}$,
where $\zeta=1+\beta\overline{\beta}$, as well as that $^{{\left(X_{(1,\beta)^{\operatorname{T}\nolimits},\gamma}\right)}\!}{\operatorname{diag}}(\alpha_{1},\alpha_{2})$ equals
$\frac{1}{\zeta^{2}}\begin{pmatrix}\alpha_{1}\zeta^{2}+(\alpha_{2}-\alpha_{1})(\gamma-1)(\overline{\gamma}-1)\beta\overline{\beta}&(\alpha_{1}-\alpha_{2})(\overline{\gamma}-1)\overline{\beta}(\gamma+\beta\overline{\beta})\\ (\alpha_{1}-\alpha_{2})(\gamma-1)\beta(\overline{\gamma}+\beta\overline{\beta})&\alpha_{2}\zeta^{2}+(\alpha_{1}-\alpha_{2})(\gamma-1)(\overline{\gamma}-1)\beta\overline{\beta}\end{pmatrix}$.
Thus, for the conclusions of the proposition to hold, we need to choose $\beta,\gamma\in\mathbb{F}_{q^{2}}$ satisfying Equation (3) together with